# Approximating the coefficients of the Bessel functions

### v2

@@ 1.5. Uniform convergence of the Bessel functions
To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.2. An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
@@ 2.1. Dunkl operators
Suppose $h\in H(\mathcal{R})$. Then, for all $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $h\mathcal{D}(\mathcal{R}(\theta))(f)h^{-1}=\mathcal{D}(\mathcal{R}(\theta))(hf)$.
@@
    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\\ \text{Re}(\left\langle ha,x\right\rangle))$.
@@ 2.2. Partitions
For $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{m})\in\Gamma$, let $|\lambda|\triangleq\sum_{i=1}^{m}\lambda_{i}$ and $\ell(\lambda)\triangleq m$. Also, we define $\Gamma_{\text{even};\,N}$ (resp. $\Gamma_{\text{even}}$) to be the set of $\lambda\in\Gamma_{N}$ (resp. $\Gamma$) such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$.
@@ 5.2. Equivariance with respect to a group action and examples
We do not consider intertwining operators in the Dunkl setting in this paper. However, we mention that the intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$. As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. As explained in [rosler_positivity], in terms of Theorem˜1.2, the intertwiner is given by
@@ 6. Leading order terms for the type A Dunkl bilinear form
However, while this proof is short, we cannot use it while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems. Therefore, we outline a new proof of Theorem˜6.1 that can be easily applied to a more general setting.
@@
(13) $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}.$
@@
Suppose $\epsilon\in\mathcal{P}$. Assume that
$M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda},$
where $c_{\epsilon\lambda}\in\mathbb{R}$ for $\lambda\in\Gamma$ are unique. Then, $c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
@@
When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}\\ p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables. We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma[|\epsilon|]}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
@@
(16) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N},\,|\lambda|=|\nu|}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}.$
@@
However, observe that $\mathcal{W}^{A}(c)[k]$ is upper-triangular with a nonzero diagonal from
Lemma˜6.16, since $c$ is not a negative integer. Thus, $\mathcal{W}^{A}(c)[k]$ is invertible, which proves (a). ∎
@@ 7. Leading order terms for the type BC Dunkl bilinear form
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
$\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that
$\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$. Consider the following statements.
@@ 8.1. Orthogonality results
For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$. Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent. We similarly have that
@@ 10. Applications
In this section, we discuss applications of the main results of this paper. Some settings that we consider include the weak convergence of the measures in ˜1.4 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
@@ 10.1. Weak convergence to the free convolution
The results of this subsection are based on the assumption of ˜1.4. Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems. In particular, we prove Corollary˜1.5. For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.7.
The following result describes the support of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$. As we mentioned earlier, [trimeche] shows that there exists a solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that is signed and is supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$. If we assume the conjecture, then we can reduce the domain of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ further, see Corollary˜10.1. We conjecture that there always exists a signed solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that has the same support as what is stated in the corollary.
Corollary 10.1.
Assume that ˜1.4 is true and that $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$. Then, for any choice of $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
@@
Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.2. Then, we have that
@@
For the sake of contradiction, assume that $\Pr_{a\sim\mu_{a_{1},a_{2}},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right]=0$. Then, there exists a sequence $\{a_{i}\}_{i\geq 1}$ of elements of $B(u,r)$ such that
$\lim_{i\rightarrow\infty}\mu_{a_{i}}^{\text{sym}}\left[B(a_{i},\frac{r}{2})\right]=0,$
since $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}\left[B(u,\frac{r}{2})\right]>0$. Let $a$ be an element of the closed ball $\overline{B}(u,r)$ that is the limit of a subsequence $\{a_{i_{j}}\}_{j\geq 1}$ of $\{a_{i}\}_{i\geq 1}$.
@@
$\lim_{j\rightarrow\infty}J_{a_{i_{j}}}^{\mathcal{R}(\theta)}(ix)=J_{a}^{\mathcal{R}(\theta)}(ix).$
Hence, it follows that the characteristic functions of $\mu^{\text{sym}}_{a_{i_{j}}}$ converge pointwise to the characteristic function of $\mu^{\text{sym}}_{a}$. By Lévy’s continuity theorem, $\mu^{\text{sym}}_{a_{i_{j}}}$ converges weakly to $\mu^{\text{sym}}_{a}$. It then follows that
$\mu^{\text{sym}}_{a}\left(B(a,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{a_{i_{j}}}\left(B(a_{i_{j}},\frac{r}{2})\right)=0,$
which is a contradiction since $a\in\text{supp}(\mu_{a}^{\text{sym}})$ by Theorem˜1.2. Thus, $u$ is in the support of $\epsilon$.
However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$. Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.2, the result follows. ∎
In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$. Afterwards, it is straightforward to deduce Corollaries˜1.5 and 1.7.
@@
Assume that ˜1.4 is true. Suppose that for $N\geq 1$, $a(N),\\ b(N)\in\mathbb{R}^{N}$. For all $N\geq 1$, assume that $\theta,\theta_{0},\theta_{1}\geq 0$.
@@
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
@@
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
@@
    for $d\geq 1$. Then, for $d\geq 1$, it is always the case that
@@
We note that the proof of Corollary˜10.2 is straightforward and that the framework of the proof has appeared previously, for example see [matrix]*Proof of Theorem 1.5.
@@
For $N\geq 1$, $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Corollary˜10.1 and is therefore exponentially decaying, see Definition˜10.15. Thus, part (C) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
@@
With this result, we are prepared to prove Corollary˜1.5.
Proof of Corollary˜1.5.
(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta(\frac{a_{i}}{\theta N})\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition. Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$. It follows from [billingsley]*Theorem 30.2 that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\right\}_{N\geq 1}$ converges in distribution to $\mu$.
(B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$. It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero. Using part (C) of Corollary˜10.2, we can then deduce that $\left\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(\frac{a_{i}}{2\theta N}\right)\right]\right\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$. Afterwards, we can follow the steps of (A). ∎
We can use the same framework to prove Corollary˜1.7.

### v3

@@ 1.5. Uniform convergence of the Bessel functions
To prove that the Bessel functions are bounded, we assume that $\theta,\theta_{0},\theta_{1}\in\mathbb{R}_{\geq 0}$ so that we can apply Theorem˜1.4. An interesting direction for future research is to generalize these arguments to whenever $\theta,\theta_{0},\theta_{1}\in\mathbb{C}$ such that the corresponding Bessel functions exist.
@@ 1.7. Applying a graded ring of operators to a graded vector field
1.8. Equivalence conditions for non-exponential holomorphic functions
An exponential formal power series is a formal power series $\exp(F_{N})$ for some $F_{N}\in\mathbb{C}[[x_{1},\ldots,x_{N}\\ ]]$. Analyzing exponential formal power series is useful since the Bessel function $J_{a}^{\mathcal{R}(\theta)}(x)$ are positive when $a,x\in\mathbb{R}^{N}$ and $\theta$ is nonnegative. Then, the Bessel generating function $\operatorname{\mathbb{E}}_{a\sim\mu}[J_{a}^{\mathcal{R}(\theta)}(x)]$, assuming that $\mu$ is supported over $\mathbb{R}^{N}$ and the integral converges, is a positive holomorphic function and can be expressed as an exponential formal power series.
We have established equivalence conditions for exponential formal power series. However, it is also straightforward to deduce analogous results for non-exponential formal power series. We describe these results in the following list.
  • In Theorems˜6.10 and 6.17, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{N}}c_{\lambda}(N)p_{\lambda}$ and replace $\exp\left(\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma}c_{\gamma}p_{\gamma}$.
  • In Theorems˜7.4 and 7.8, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}$ and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$.
  • In Theorems˜8.21 and 8.26, we set $F_{N}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma_{\text{even};\,N}}(c_{\lambda}(N)+ed_{\lambda}(N))p_{\lambda}$, replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}$, and replace $\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}\right)$ with $\sum_{\gamma\in\Gamma_{\text{even}}}d_{\gamma}p_{\gamma}$.
@@ 1.9. Paper organization
1.10. Acknowledgments
The author would like to thank Vadim Gorin for giving comments on the paper.
@@ 2.1. Dunkl operators
Suppose $h\in H(\mathcal{R})$. Then, for all $f\in\mathbb{C}[x_{1},\ldots,x_{N}]$, $h\mathcal{D}(\mathcal{R}(\theta))\\ (f)h^{-1}=\mathcal{D}(\mathcal{R}(\theta))(hf)$.
@@
    If $\text{Re}(\theta(r))\geq 0$ for all $r\in\mathcal{R}$, then $|E_{a}^{\mathcal{R}(\theta)}(x)|\leq\sqrt{|H(\mathcal{R})|}\exp(\max_{h\in H(\mathcal{R})}\text{Re}(\\ \left\langle ha,x\right\rangle))$.
@@ 2.2. Partitions
For $\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{m})\in\Gamma$, let $|\lambda|\triangleq\sum_{i=1}^{m}\lambda_{i}$ and $\ell(\lambda)\triangleq m$. Also, we define $\Gamma_{\text{even};\,N}$ (resp. $\Gamma_{\text{even}}$) to be the set of $\lambda\in\Gamma_{N}$ (resp. $\Gamma$) such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$. Similarly, define $\mathcal{P}_{\text{even};\,N}$ to be the set of $\lambda\in\mathcal{P}_{N}$ such that $\lambda_{i}$ is even for all $i\in[\ell(\lambda)]$.
@@
Additionally, define the unique coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$ such that
$M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}.$
Remark 2.13.
When we write $M_{\epsilon}=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}$, we mean that $M_{\epsilon}(x)=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x)$, where $x=(x_{i})_{i\geq 1}$ consists of an infinite number of variables. We could equivalently state that $M_{\epsilon}(x_{1},\ldots,x_{N})=\sum_{\lambda\in\Gamma}c_{\epsilon\lambda}p_{\lambda}(x_{1},\ldots,x_{N})$ for all $N\geq 1$.
@@ 5.2. Equivariance with respect to a group action and examples
We do not consider intertwining operators in the Dunkl setting in this paper. However, we mention that the intertwining operator for $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ is well-studied, where $\partial\in\mathcal{T}$ satisfies $\partial(x_{i})=\partial_{i}$ for $i\in[N]$; equivalently, $\partial\triangleq\mathcal{D}(\mathcal{R}(0))$. As an example of an application of Theorem˜4.5, the theorem implies the well-known result that $(\mathcal{D}(\mathcal{R}(\theta)),\partial)$ intertwines with a unique operator if and only if $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. As explained in [rosler_positivity], in terms of Theorem˜1.4, the intertwiner is given by
@@ 6. Leading order terms for the type A Dunkl bilinear form
Although this proof is short, we cannot use it while proving analogous results for the $BC^{N}(\theta_{0},\theta_{1})$ and $D^{N}(\theta)$ root systems. Therefore, we outline a new proof of Theorem˜6.1 that can be easily applied to a more general setting.
@@
(13) $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[|\nu|]}(\mathcal{M}^{A}_{\nu\lambda}+o_{N}(1))\frac{[p_{\lambda}]F_{N}}{(\theta N)^{\ell(\lambda)}}.$
@@
Suppose $\epsilon\in\mathcal{P}$. Then, $c_{\epsilon,\,(|\epsilon|)}=(-1)^{\ell(\epsilon)-1}\frac{|\epsilon|!}{\pi(\epsilon)\ell(\epsilon)}$.
@@
See Subsection˜2.2 for the definition of $c_{\epsilon,\,(|\epsilon|)}$.
@@
(16) $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{A^{N-1}(\theta)}=\lim_{N\rightarrow\infty}\sum_{\lambda\in\Gamma_{N}[|\nu|]}(\mathcal{W}^{A}(c)_{\nu\lambda}+o_{N}(1))[p_{\lambda}]F_{N}.$
@@
However, observe that $\mathcal{W}^{A}(c)[k]$ is upper-triangular with a nonzero diagonal from Lemma 6.16, since $c$ is not a negative integer. Thus, $\mathcal{W}^{A}(c)[k]$ is invertible, which proves (a). ∎
@@ 7. Leading order terms for the type BC Dunkl bilinear form
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta_{0}N|=\infty$ and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Consider the following statements.
@@
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta_{0}N=c_{0}\in\mathbb{C}$ and $\lim_{N\rightarrow\infty}\theta_{1}=c_{1}\in\mathbb{C}$. Consider the following statements.
@@ 8.1. Orthogonality results
For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$. Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B^{N}(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent. We similarly have that
@@ 10. Applications
In this section, we discuss applications of the main results of this paper. Some settings that we consider include the weak convergence of the measures in Conjecture˜1.6 to the free convolution and the uniform convergence of the Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime and for Vershik-Kerov sequences.
@@ 10.1. Weak convergence to the free convolution
The results of this subsection are based on the assumption of Conjecture˜1.6. Afterwards, the goal is to show that if $a_{1}$ and $a_{2}$ converge weakly to some distributions, then $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ also converges weakly to some distribution that can be described in terms of the free convolution, if we are working with the type A and D root systems. In particular, we prove Corollary˜1.7. For the type BC root system, we show weak converge to the rectangular free convolution, see Corollary˜1.8.
The following result describes the support of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$. As we mentioned earlier, [trimeche] shows that there exists a solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that is signed and is supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$. If we assume the conjecture, then we can reduce the domain of $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ further, see Lemma˜10.1. We conjecture that there always exists a signed solution for $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ that has the same support as what is stated in the corollary.
Lemma 10.1.
Assume that $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$ and $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, every element of $\text{supp}(\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}})$ can be expressed as $u_{1}+u_{2}$, where $u_{i}$ is in the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$.
@@
Consider the measures $\mu_{a}^{\text{sym}}$ for $a\in\mathbb{R}^{N}$ as they are defined in Theorem˜1.4. Then, we have that
@@
For the sake of contradiction, assume that $\Pr_{a\sim\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)},\,\epsilon^{\prime}\sim\mu_{a}^{\text{sym}}}\left[a\in B(u,\frac{r}{2}),\epsilon^{\prime}\in B(a,\frac{r}{2})\right]=0$. Then, there exists a sequence $\{z_{i}\}_{i\geq 1}$ of elements of $B(u,r)$ such that
$\lim_{i\rightarrow\infty}\mu_{z_{i}}^{\text{sym}}\left[B(z_{i},\frac{r}{2})\right]=0,$
since $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}\left[B(u,\frac{r}{2})\right]>0$. Let $z$ be an element of the closed ball $\overline{B}(u,r)$ that is the limit of a subsequence $\{z_{i_{j}}\}_{j\geq 1}$ of $\{z_{i}\}_{i\geq 1}$.
@@
$\lim_{j\rightarrow\infty}J_{z_{i_{j}}}^{\mathcal{R}(\theta)}(ix)=J_{z}^{\mathcal{R}(\theta)}(ix).$
Hence, it follows that the characteristic functions of $\mu^{\text{sym}}_{z_{i_{j}}}$ converge pointwise to the characteristic function of $\mu^{\text{sym}}_{z}$. By Lévy’s continuity theorem, $\mu^{\text{sym}}_{z_{i_{j}}}$ converges weakly to $\mu^{\text{sym}}_{z}$. It then follows that
$\mu^{\text{sym}}_{z}\left(B(z,\frac{r}{4})\right)=\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z,\frac{r}{4})\right)\leq\lim_{j\rightarrow\infty}\mu^{\text{sym}}_{z_{i_{j}}}\left(B(z_{i_{j}},\frac{r}{2})\right)=0,$
which is a contradiction since $z\in\text{supp}(\mu_{z}^{\text{sym}})$ by Theorem˜1.4. Thus, $u$ is in the support of $\epsilon$.
However, the distribution of $\epsilon$ is the same as that of $\epsilon^{1}+\epsilon^{2}$, so $u$ is in the support of $\epsilon^{1}+\epsilon^{2}$. Since the support of $\epsilon^{i}$ is a subset of the convex hull of $H(\mathcal{R})a_{i}$ for $i\in\{1,2\}$ by Theorem˜1.4, the result follows. ∎
In the following corollary, we compute the convergence of the measures $\mu_{a,b}^{\mathcal{R}(\theta)}$ in terms of moments after assuming that $a$ and $b$ converge in terms of moments as $N\rightarrow\infty$. Afterwards, it is straightforward to deduce Corollaries˜1.7 and 1.8.
@@
Suppose that for $N\geq 1$, $a(N),b(N)\in\mathbb{R}^{N}$. For all $N\geq 1$, assume that $\theta,\theta_{0},\theta_{1}\geq 0$.
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
@@
    for $d\geq 1$. Assume that for $N\geq 1$, $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, for $d\geq 1$,
@@
We note that the proof of Corollary˜10.2 is straightforward and that its framework has appeared previously, for example see [matrix]*Proof of Theorem 1.5. The main contribution of the proof is the application of Corollary˜6.11.
@@
For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15. Thus, part (E) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
@@
With this result, we are prepared to prove Corollary˜1.7.
@@
(A): By part (A) of Corollary˜10.2, $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\\ \Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. Since $\mu_{a}$ and $\mu_{b}$ are compactly supported, $\mu$ is also compactly supported by definition. Then, by Carleman’s condition, we can deduce that the measure is the only measure over $\mathbb{R}$ with moments equal to the moments of $\mu$. It follows from [billingsley]*Theorem 30.2 that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in distribution to $\mu$.
(B): Note that the odd moments of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ are always zero, so therefore the odd free cumulants of the measure are also always zero, and similarly for $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$. It follows that the odd free cumulants of $\tilde{\mu}$ are always zero and therefore that the odd moments of $\tilde{\mu}$ are always zero. Using part (C) of Corollary˜10.2, we can then deduce that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\tilde{\mu}$. Afterwards, we can follow the steps of (A). ∎
We can use the same framework to prove Corollary˜1.8.
Proof of Corollary˜1.8.